Does 3^5+3^5+3^5= 3^6. Please explain.
Answer by bucky(2189)

The answer to this question is "Yes".
.
3^5 equals 3*3*3*3*3 = 243
.
So:
.
3^5+3^5+3^5 = 243 + 243 +243 = 729
.
But also:
.
3^6 = 3*3*3*3*3*3 = 729
.
The reason for this may be easier to see if you start with the original expression:
.
3^5 + 3^5 + 3^5
.
Substitute X for 3^5 and this expression becomes:
.
X + X + X
.
But this expression adds up to be 3X.
.
Now replace the X with its equal 3^5 and you can see that you now have
.
3*3^5
.
The multiplier 3 has an implied exponent of 1. So this expression is equivalent to:
.
3^1 * 3^5
.
And, since both multipliers have the base 3, you multiply them by adding the exponents:
.
3^1 * 3^5 = 3^(1+5) = 3^6
.
Hope this helps you to see why:
.
3^5+3^5+3^5 = 3*3^5 = 3^6
.